OpenStudy (anonymous):

Find the equation of the secant line containing (1, f(1)) and (6, (f(6)) f(x)=x^2-2x Write the coefficients of the equation in the form of a simple fraction

OpenStudy (anonymous):

The secant line is the average slope, use the standard equation of a slope. \[\frac{ y1-y0 }{ x1-x0 }\] You have two points (x=1,y=f(1)) and (x=6,y+f(6)) solve for f(1) and f(6) and use the slope equation.

OpenStudy (anonymous):

f(1)=-2 and f(6)=24 right?

OpenStudy (anonymous):

\[f(1)=1^{2}-2(1)= -1\] \[f(6)=6^{2}-2(6)=24\]

OpenStudy (anonymous):

where do you plug those values in?

OpenStudy (anonymous):

Think of this "f(x)" as a function. You give me an "x" value and I will tell you what "y" is. This is because "y" is DEPENDENT on what the "x" value is. "x" is the INDEPENDENT variable. Plug the value given, lets say 6 in the case of "f(6)" into your equation "f(x)" and f(x)=x^(2)-2x

OpenStudy (anonymous):

Then take those values as being your two "y" points. Remember which one is y1 nad which is y0 and plug that into the slope equation.

OpenStudy (anonymous):

got it! thank youuu

OpenStudy (anonymous):

No problemo